### v1

# exampleframelinecolor=gray!30, linewidth=0.5pt, backgroundcolor=gray!5, roundcorner=2pt, innertopmargin=8pt, innerbottommargin=8pt, innerleftmargin=8pt, innerrightmargin=8pt, skipabove=skipbelow=

Abstract
Large Language Models (LLMs) rely on safety alignment to produce socially acceptable responses. This is typically achieved through instruction tuning and reinforcement learning from human feedback. However, this alignment is known to be brittle: further fine-tuning, even on benign or lightly contaminated data, can degrade safety and reintroduce harmful behaviors. A growing body of work suggests that alignment may correspond to identifiable geometric directions in weight space, forming subspaces that could, in principle, be isolated or preserved to defend against misalignment. In this work, we conduct a comprehensive empirical study of this geometric perspective. We examine whether safety-relevant behavior is concentrated in specific subspaces, whether it can be separated from general-purpose learning, and whether harmfulness arises from distinguishable patterns in internal representations. Across both parameter and activation space, our findings are consistent: subspaces that amplify safe behaviors also amplify unsafe ones, and prompts with different safety implications activate overlapping representations. We find no evidence of a subspace that selectively governs safety. These results challenge the assumption that alignment is geometrically localized. Rather than residing in distinct directions, safety appears to emerge from entangled, high-impact components of the model’s broader learning dynamics. This suggests that subspace-based defenses may face fundamental limitations and underscores the need for alternative strategies to preserve alignment under continued training. We corroborate these findings through multiple experiments on five open-source LLMs. Our code is publicly available at: https://github.com/CERT-Lab/safety-subspaces.
1 Introduction
Large Language Models (LLMs) have shown strong performance across a wide range of general-purpose tasks, including complex reasoning and problem solving [43, 1, 51, 49, 58, 50]. To ensure these models behave responsibly and align with human values, they undergo an additional process of security alignment. Despite known jailbreak methods that can bypass safeguards, aligned models are generally considered significantly safer than their base versions [37, 55, 44]. This alignment is typically achieved through supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF), enabling models only improve response quality, and also refuse harmful or inappropriate prompts. A growing line of research focuses on the weight difference between the base and aligned models, commonly referred to as the alignment matrix, which captures the transition from unaligned to aligned behavior. This difference has been used to interpret alignment mechanisms and develop defenses against adversarial attacks [3, 18, 57, 10, 30, 25, 28].
However, this alignment is fragile. Since safety is encoded in the model’s weights, any modification, such as further fine-tuning (FT), can compromise it. While FT adapts models to new tasks by learning update directions, it offers no guarantee that safety is preserved. This exposes a deeper attack surface beyond prompt engineering: an adversary could insert a small number of malicious samples into a training set to subvert alignment [59, 60, 4, 64]. Recent work shows that even benign FT, low-rank adaptation, or pruning can degrade a model’s safety profile [29, 42, 15, 16, 54]. Preserving alignment under continued training is therefore both a practical concern and a theoretically challenging problem. Given this vulnerability, a natural question arises: Does there exist any subspace, whether in weight space or activation space, that uniquely encodes safety alignment? If safety is a distinct and structured property of the model, then updates or representations affecting it might consistently concentrate in identifiable geometric regions. This motivates a broader question: can we isolate or characterize safety-relevant subspaces that amplify aligned behavior or suppress harmful outputs?
To explore this question, we conduct four experiments probing the geometry of safety-related behavior across model weights and internal representations. We begin by analyzing FT updates from purely useful and harmful datasets, projecting them into subspaces derived from the alignment matrix to test whether safety correlates with energy or behavioral impact. Next, we examine contaminated FT, where a small fraction of harmful samples is mixed into a benign dataset. By projecting updates into the orthogonal complement of alignment-derived subspaces, we test whether harmful components can be selectively removed (see Figure 1). In the third experiment, we directly compare the dominant subspaces of useful, harmful, and alignment updates to assess whether safety-altering updates share consistent structure. Finally, we inspect the representation space, comparing internal activations from useful and harmful prompts to ask whether safety-related inputs occupy distinct subspaces, even when weight updates do not.
Across all experiments, we observe a consistent and surprising result: no subspace, whether defined by alignment directions, update energies, or input representations, captures safety-specific behavior in isolation. While certain subspaces, such as the top alignment directions, are behaviorally impactful, they amplify both helpful and harmful behaviors equally, reflecting general update sensitivity rather than alignment. Similarly, activations from harmful and helpful prompts occupy overlapping regions of representation space, offering no evidence for distinct "safety activation" geometry. These findings point to a fundamental limitation of subspace-based alignment strategies. If safe and unsafe behaviors cannot be cleanly separated geometrically, then projection- or filtering-based defenses are unlikely to suppress harmfulness without incurring equivalent losses in utility. Our key contributions are:
  • We show that subspaces derived from alignment updates are not safety-specific; they amplify both helpful and harmful behaviors equally, reflecting general update sensitivity rather than alignment.
  • We find that orthogonal projection intended to filter harmful updates leads to proportional losses in utility, suggesting no selective geometric separation between safe and unsafe behavior.
  • We demonstrate that harmful and aligned updates do not share a consistent subspace, and that harmful prompts do not activate distinct regions of representation space.
  • Through consistent results across five open-source LLMs evaluated in multiple experiments, we challenge the view that safety alignment is geometrically localized and reveal fundamental limitations of subspace-based defenses.
2 Preliminaries
Notation.
Let $\mathbf{W}_{0}$ denote the parameters of the base model, and let $\mathbf{W}_{\mathrm{A}}$ represent the parameters of the aligned instruction-tuned model. We further fine-tune the aligned model on a task-specific dataset $\mathcal{D}j$, where $j\in\{\text{Useful},\text{Harmful},\text{Contaminated}\}$, resulting in parameters $\mathbf{W}_{\mathrm{FT},j}$. We decompose the total parameter update as the sum of two components:
[…]
Importance of Alignment Directions ($\Delta_{\mathrm{A}}$).
Alignment training typically emphasizes behavioral properties such as harmlessness, helpfulness, and honesty. Empirical studies [37, 18, 10] suggest that the alignment update $\Delta_{\mathrm{A}}$ encodes directions in parameter space that are strongly correlated with these safety attributes. Our goal is to systematically control the extent to which the subsequent task-specific update $\Delta_{\mathrm{T}}^{j}$ interacts with these alignment directions.
Figure 1: The base model $W_{o}$ is instruction-tuned to produce the aligned model $W_{A}$. Step 1: The difference $\Delta_{A}=W_{A}-W_{o}$ defines a safety direction, from which projection matrices $P_{k}$ (top-K subspace) and $P_{k}^{\perp}$ (orthogonal subspace) are derived. $W_{A}$ is then fine-tuned on three datasets: helpful, harmful, and contaminated, to yield $W_{\text{useful}}$, $W_{\text{harmful}}$, and $W_{\text{contaminated}}$, with updates $\Delta_{t_{j}}$. Step 2: Project $\Delta_{t_{j}}$ using $P_{k}$ and $P_{k}^{\perp}$, and add back to $W_{A}$ to obtain projected models for evaluation. In addition, SVD is performed on the task-specific updates and MSO is computed between the top-K singular vectors.
Constructing the Alignment Subspace.
To formalize this notion, we begin by constructing the alignment subspace. Each tensor in the alignment update $\Delta_{\mathrm{A}}$ is first reshaped into a matrix (flattened if needed) $V_{\mathrm{A}}\in\mathbb{R}^{M\times N}$. We perform a thin singular value decomposition (SVD) of the form $V_{\mathrm{A}}=U\Sigma V^{\top}$, which reveals the principal directions of parameter change [12, 36], ranked by their contribution to the Frobenius norm. The top $k$ (Top-K) right singular vectors in $V$ are then selected to define the alignment subspace:
$\mathcal{S}_{k}:=\operatorname{span}(U_{k}),\quad U_{k}\in\mathbb{R}^{M\times k},\quad k\leq\operatorname{rank}(V_{\mathrm{A}}).$ (3)
Intuitively, $\mathcal{S}_{k}$ captures the $k$ most significant directions of parameter shifts resulting from alignment training. The alignment subspace naturally induces projection operators:
$P_{k}:=U_{k}U_{k}^{\!\top},\quad P_{k}^{\perp}:=I-P_{k},$ (4)
where $P_{k}$ projects a matrix onto the alignment subspace, and $P_{k}^{\perp}$ onto its orthogonal complement.
Projection Schemes.
Given a fractional rank hyperparameter $\varrho\in(0,1]$, we determine $k=\lfloor\varrho\cdot\min(M,N)\rfloor$ and apply one of two projection-based update schemes to the task-specific update:
[…]
Equation 5 retains the update components that align with the alignment directions, while Equation 6 removes this aligned component, retaining only the update orthogonal to the alignment subspace.
Control Experiments.
To further assess the specificity and effectiveness of the alignment subspace, we introduce two control experiments:
  • Random-K: Instead of using the top-$k$ singular vectors from the SVD of $V_{\mathrm{A}}$, we randomly sample $k$ singular vectors from the full set to construct a randomized alignment subspace.
  • Random: We replace $V_{\mathrm{A}}$ with a random matrix of the same dimensions, perform its SVD, and use the top-$k$ singular vectors to define a synthetic alignment subspace.
Energy-Kept Ratio.
We introduce the fractional energy metric to quantify the extent of overlap between the task update and the alignment subspace:
$\mathcal{E}_{k}(\Delta_{\mathrm{T}}^{j}):=\frac{\|P_{k}\Delta_{\mathrm{T}}^{j}\|_{F}^{2}}{\|\Delta_{\mathrm{T}}^{j}\|_{F}^{2}},\quad\mathcal{E}_{k}^{\perp}(\Delta_{\mathrm{T}}^{j})=1-\mathcal{E}_{k}(\Delta_{\mathrm{T}}^{j}).$ (7)
Mode Subspace Overlap (MSO).
Let $\mathbf{V}\in\mathbb{R}^{d\times n_{V}}$ and $\mathbf{W}\in\mathbb{R}^{d\times n_{W}}$ be two matrices with a shared ambient dimension $d$ but possibly different column counts. We extract their principal directions using thin SVD:
$\mathbf{V}=U_{V}\Sigma_{V}V_{V}^{\top},\quad\mathbf{W}=U_{W}\Sigma_{W}V_{W}^{\top}.$ (8)
For a chosen energy-retention fraction $\eta\in(0,1]$, we select the smallest $k_{V}$ and $k_{W}$ such that the top $k_{V}$ (resp. $k_{W}$) left singular vectors capture at least an $\eta$-fraction of $\|\Sigma_{V}\|_{F}^{2}$ (resp. $\|\Sigma_{W}\|_{F}^{2}$). This yields orthonormal bases $Q_{V}\in\mathbb{R}^{d\times k_{V}}$ and $Q_{W}\in\mathbb{R}^{d\times k_{W}}$. The overlap matrix is then defined as:
$S=Q_{V}^{\top}Q_{W}\in\mathbb{R}^{k_{V}\times k_{W}}.$ (9)
To quantify the similarity between these $\eta$-energy subspaces, we use MSO metric:
$\mathrm{MSO}(\mathbf{V},\mathbf{W};\eta)=\frac{\lVert S\rVert_{F}^{2}}{\min(k_{V},k_{W})},\quad 0\leq\mathrm{MSO}\leq 1.$ (10)
Intuitively, $\mathrm{MSO}(\mathbf{V},\mathbf{W};\eta)$ measures the overlap between the top-$\eta$ energy components of $\mathbf{V}$ and $\mathbf{W}$: it equals 0 for orthogonal subspaces and 1 for identical spans. As a baseline, the expected overlap between random subspaces of dimensions $k_{V}$ and $k_{W}$ in $\mathbb{R}^{d}$ is given analytically by:
$\mathbb{E}[\mathrm{overlap}]=\frac{\max(k_{V},k_{W})}{d}.$ (11)
Models Used.
Throughout our work, we evaluate both base and instruction-aligned versions of several open-source LLMs. For example, we consider Qwen-2.5 3B (base) alongside its aligned variant, Qwen-2.5 3B Instruct. We report results for base and aligned versions of five models: LLaMA 3.2 1B [11], LLaMA 2 7B [51], Qwen-2.5 1B [58], Qwen-2.5 3B, and Qwen-2.5 7B.
3 Do Alignment Subspaces Encode Safety?
Figure 2: Parallel projection-based update schemes across varying SVD fractions. We report the energy-kept ratio for models fine-tuned on Full Useful and Full Harmful data, utility for models fine-tuned on Full Useful, and harmfulness for models fine-tuned on Full Harmful.
A central question in understanding safety alignment is whether specific directions in weight space, such as those defined by the difference between a base model and its RLHF-aligned counterpart, encode information unique to safety. If this is the case, then constraining FT updates to lie within these subspaces could offer a principled way to guard against harmful optimization. We begin our investigation by examining whether task-specific FT updates align differently with the top directions of the alignment matrix, depending on whether the task is helpful or harmful.
Experimental Setup.
We fine-tune an aligned instruction-tuned model on two distinct datasets. The first is a 20K subset of MetaMathQA [63], a benchmark of math word problems representing a useful task without safety concerns. The second is a 4K unsafe subset of BeaverTails [26], a synthetic dataset of harmful instruction–response pairs designed to elicit unsafe behavior. We denote the resulting weight updates as $\Delta_{\mathrm{T}}^{\text{Useful}}$ and $\Delta_{\mathrm{T}}^{\text{Harmful}}$ respectively. To quantify behavioral effects, we evaluate harmfulness using the AdvBench dataset [68], with GPT-4o-mini [24] scoring each response from 1 (least harmful) to 5 (most harmful); the final score is the average across samples. Utility is measured by accuracy on the GSM8k test set [8], using final answer correctness. We compute these metrics, energy-kept ratio, utility, and harmfulness, for the projected models $\mathbf{W}_{\text{parallel}}$ and $\mathbf{W}_{\text{orthogonal}}$, as well as for the base, aligned, fine-tuned, and control models.
Results: Energy Is Uniform Across Subspaces, Performance Is Not.
As shown in Figure 2, the fraction of energy retained in projected updates increases linearly with subspace rank and is consistent across all three subspace types. This pattern holds for both helpful and harmful updates. There is no evidence that update energy is preferentially concentrated in the top directions of $\Delta_{A}$ for safe vs unsafe FT. This suggests that if a "safety subspace" exists, it is not captured simply by energetic alignment with the dominant directions of $\Delta_{A}$. However, while energy is evenly distributed, behavioral impact is not. We can observe that projecting $\Delta_{\mathrm{T}}^{\text{Useful}}$ onto the top-$k$ directions consistently improves utility relative to random projections with equal energy, in Figure 2 and Table 1. Similarly, projecting $\Delta_{\mathrm{T}}^{\text{Harmful}}$ onto the same directions increases harmfulness. Thus, the top singular directions of $\Delta_{A}$ are not uniquely aligned with safety, but they are generally potent. Updates along these directions are more effective, whether the goal is to enhance utility or to elicit harmful behavior. We present results on all models in Table 3 (Appendix B).
Table 1: Parallel projection-based update schemes across varying SVD fractions. We report the utility for models fine-tuned on Full Useful data, and harmfulness for models fine-tuned on Full Harmful.
Model Method Utility ($\uparrow$) Harmful Score ($\downarrow$)
Aligned SVD Fractions FT Aligned SVD Fractions FT
0.01 0.25 0.50 0.75 0.99 0.01 0.25 0.50 0.75 0.99
Qwen-2.5 1.5B Top-K $0.47$ $0.50$ $0.53$ $0.55$ $0.57$ $0.58$ $0.61$ $1.55$ $1.62$ $1.80$ $1.92$ $1.90$ $1.97$ $2.09$
Random-K $0.47$ $0.49$ $0.50$ $0.53$ $0.56$ $0.58$ $0.61$ $1.55$ $1.55$ $1.66$ $1.78$ $1.92$ $2.00$ $2.09$
Random $0.47$ $0.49$ $0.50$ $0.53$ $0.53$ $0.56$ $0.61$ $1.55$ $1.56$ $1.65$ $1.74$ $1.83$ $1.95$ $2.09$
Llama-3.2 1B Top-K $0.13$ $0.14$ $0.21$ $0.25$ $0.30$ $0.34$ $0.36$ $2.80$ $2.89$ $3.29$ $3.51$ $3.66$ $3.84$ $4.07$
Random-K $0.13$ $0.13$ $0.16$ $0.23$ $0.29$ $0.34$ $0.36$ $2.80$ $2.83$ $3.11$ $3.37$ $3.55$ $3.84$ $4.07$
Random $0.13$ $0.13$ $0.17$ $0.22$ $0.29$ $0.34$ $0.36$ $2.80$ $2.81$ $3.05$ $3.34$ $3.56$ $3.83$ $4.07$
Implications: Alignment Directions Reflect Update Sensitivity, Not Safety.
This symmetry across tasks is important. The fact that top-$k$ directions amplify both helpful and harmful behavior equally suggests they do not encode alignment directly. Instead, they appear to represent axes of general parameter sensitivity, directions where updates tend to induce large changes in model behavior. In this sense, $\Delta_{A}$ captures a general learning geometry: directions that are especially effective for optimization, not inherently safe. We draw three key takeaways. First, neither helpful nor harmful updates preferentially align with the top subspaces of $\Delta_{A}$ in terms of energy. Second, those same subspaces are more behaviorally expressive, enhancing both utility and harmfulness depending on the task. Third, this challenges the notion that $\Delta_{A}$ encodes safety-specific information. Its dominant directions support effective learning broadly, without guiding its ethical character. Thus, using $\Delta_{A}$ to constrain updates may regulate the magnitude of behavior change, but not its direction or valence.
4 Can Harmful Subspaces Be Removed?
Having analyzed helpful and harmful updates in isolation, we now consider a more realistic scenario: contaminated FT. This involves adding a small fraction of harmful examples to an otherwise benign dataset, producing updates that blend both signals. Contaminated data is particularly dangerous because it can degrade alignment without obvious signs. Prior work shows that even limited contamination can erode safety, causing models to revert to unsafe behaviors. While earlier experiments identified expressive subspaces, we now ask the reverse: can we remove harmful components from an update? We test whether filtering specific subspaces, particularly those aligned with the dominant directions of the alignment matrix, can reduce harmfulness while preserving utility.
Experimental Setup.
We construct a contaminated dataset by mixing 20% harmful data from BeaverTails with 80% of the 20K MetaMathQA subset. FT on this mixture yields a single contaminated update, $\Delta_{T}$. To suppress harmful behavior, we apply the orthogonal projection strategy from Section 2, removing components along the top-$k$ alignment directions. Specifically, we compute $\tilde{\Delta}_{T}=P_{k}^{\perp}\Delta_{T}$, where $P_{k}^{\perp}$ projects onto the complement of the alignment subspace. We evaluate the resulting models on GSM8K (utility) and AdvBench (harmfulness). Our goal is to test whether removing alignment-aligned components can reduce harmfulness while preserving task performance.
Results: Utility And Harmfulness Drop Together.
Figure 3 shows the effects of orthogonal projection on retained energy, utility, and harmfulness. As $k$ increases, implying more of the update is removed, the retained energy declines steadily across all projection types (random, top-$k$, and random-$k$). Utility and harmfulness scores (Figure 3, Table 2) follow a similar downward trend. However, the rate of decline differs by projection strategy. Removing top-$k$ alignment components reduces utility more sharply than random projections. At the same time, harmfulness decreases at a similar rate, indicating no selective suppression of harmful behavior. In effect, safety improvements come at a proportional cost to task performance, with no clear advantage in targeting the alignment subspace. We present results on all models in Table 4 (Appendix C).
Figure 3: Parallel projection-based update schemes across varying SVD fractions. We report the energy-kept ratio for models fine-tuned on Full Useful, Full Harmful and Contaminated data; and utility and harmfulness for models fine-tuned on Contaminated.
Table 2: Parallel projection-based update schemes across varying SVD fractions. We report the utility and harmfulness for models fine-tuned on Contaminated data.
Model Method Utility ($\uparrow$) Harmful Score ($\downarrow$)
Aligned SVD Fractions FT Aligned SVD Fractions FT
0.01 0.25 0.50 0.75 0.99 0.01 0.25 0.50 0.75 0.99
Qwen-2.5 1.5B Top-K $0.47$ $0.50$ $0.53$ $0.55$ $0.57$ $0.58$ $0.60$ $1.55$ $1.58$ $1.65$ $1.80$ $1.91$ $1.92$ $2.16$
Random-K $0.47$ $0.49$ $0.52$ $0.53$ $0.55$ $0.55$ $0.60$ $1.55$ $1.56$ $1.62$ $1.63$ $1.87$ $1.92$ $2.16$
Random $0.47$ $0.49$ $0.50$ $0.52$ $0.52$ $0.54$ $0.61$ $1.55$ $1.58$ $1.64$ $1.68$ $1.74$ $1.92$ $2.16$
Qwen-2.5 3B Top-K $0.61$ $0.63$ $0.64$ $0.65$ $0.68$ $0.69$ $0.73$ $1.47$ $1.49$ $1.58$ $1.69$ $1.76$ $1.83$ $1.99$
Random-K $0.61$ $0.62$ $0.64$ $0.64$ $0.66$ $0.69$ $0.73$ $1.47$ $1.45$ $1.55$ $1.62$ $1.65$ $1.91$ $1.99$
Random $0.61$ $0.62$ $0.63$ $0.64$ $0.65$ $0.68$ $0.73$ $1.47$ $1.45$ $1.50$ $1.57$ $1.75$ $1.83$ $1.99$
Qwen-2.5 7B Top-K $0.74$ $0.74$ $0.75$ $0.75$ $0.75$ $0.78$ $0.81$ $1.30$ $1.31$ $1.56$ $1.60$ $1.68$ $1.67$ $1.96$
Random-K $0.74$ $0.74$ $0.75$ $0.76$ $0.75$ $0.78$ $0.81$ $1.30$ $1.35$ $1.41$ $1.46$ $1.59$ $1.67$ $1.96$
Random $0.74$ $0.74$ $0.75$ $0.75$ $0.75$ $0.78$ $0.81$ $1.30$ $1.34$ $1.40$ $1.48$ $1.56$ $1.63$ $1.96$
Implications: No Selective Removal Is Possible.
These results indicate that the alignment subspace does not uniquely encode safety or harmfulness but rather captures directions broadly important for learning. Removing these directions degrades both utility and harmfulness at similar rates. If harmful behavior were confined to distinct subspaces, we would expect a steeper drop in harmfulness than utility, yet this is not observed. Even if safety-relevant directions exist, they are not recoverable from the alignment matrix alone, especially under contamination. The update blends helpful and harmful objectives, making its projection agnostic to intent. As a result, orthogonal projection fails to selectively suppress harmful behavior. Subspace filtering based on alignment directions imposes a strict tradeoff: gains in safety come with proportional utility loss. This challenges the effectiveness of subspace-based defenses under contaminated FT.
5 Are Safety Weight Subspaces Distinct?
Figure 4: Mode Subspace Overlap (MSO) at the 70- and 85- percentile layers for pairwise comparisons of the dominant subspaces from Harmful fine-tuned (H), Instruction-tuned (I), and Base (B) models.
A natural question is whether a dedicated region of parameter space, what we might call a safety subspace, captures safety-specific behavior. Such a subspace should meet two criteria: (i) safety-relevant updates, whether from alignment or harmful FT, should lie significantly within it; and (ii) task-specific updates unrelated to safety should have minimal overlap, with projections onto the subspace leaving model safety unchanged. Crucially, these properties must generalize across tasks and datasets to rule out dataset-specific artifacts. Our earlier results argue against the top subspaces of the alignment matrix meeting these criteria. These directions are highly sensitive to any update, helpful or harmful, but do not isolate safety. Still, it remains open whether some other set of directions, possibly outside the alignment subspace, could fulfill this role. To explore this, we directly compare the dominant subspaces of different update types.
Experimental Setup.
We compare the principal subspaces of 3 updates: the alignment update $\Delta_{A}$ (from base to aligned model), the harmful FT update $\Delta_{\mathrm{T}}^{\text{Harmful}}$ (trained on BeaverTails), and the useful update $\Delta_{\mathrm{T}}^{\text{Useful}}$ (trained on a 20K subset of MetaMathQA). Notably, the negated alignment update, $-\Delta_{A}$, reverses alignment by pushing the model back toward its unaligned base state, effectively acting as a harmful update and serving as a useful reference. For a given energy threshold $\eta\in(0,1]$, we compute $\mathrm{MSO}(\cdot,\cdot;\eta)$ (Section 2 for three pairs: (i) $\big{(}\Delta_{\mathrm{T}}^{\text{Useful}},\Delta_{\mathrm{T}}^{\text{Harmful}}\big{)}$, to assess whether helpful and harmful FT affect similar subspaces; (ii) $\big{(}\Delta_{\mathrm{T}}^{\text{Useful}},-\Delta_{A}\big{)}$, to test alignment between helpful updates and reversed alignment; and (iii) $\big{(}\Delta_{\mathrm{T}}^{\text{Harmful}},-\Delta_{A}\big{)}$, to compare two harmful directions. We sweep over $\eta$, with small values isolating high-energy directions and larger values approaching full-rank overlap. We include the random-subspace baseline $\max(k_{V},k_{W})/d$; values above this baseline indicate significant geometric alignment, while values near it suggest chance-level overlap.
Results: Representations Overlap Across Tasks.
Figure 4 shows the pairwise overlap between the dominant subspaces (top-$k$ directions) of each update. All pairs exhibit greater overlap than random baselines, indicating shared structure. However, the strongest overlap is between the useful and harmful updates, not between alignment and harmful updates, as one might expect if safety were a shared component. This is a key finding. If a safety subspace existed, it would likely appear in the shared directions between alignment and harmful updates, which affect safety in opposite ways. This lack of substantial overlap suggests that no consistent, linear safety-specific subspace exists.
Implications: Shared Subspaces Drive Behavior, Not Safety.
Together with earlier results, these findings suggest that safety-relevant updates do not lie in a well-defined or isolatable subspace. Instead, both alignment and harmfulness operate over complex, task-dependent, and likely non-linear directions. The high overlap between harmful and helpful update subspaces supports our earlier hypothesis: these directions form a general learning subspace, expressive across tasks but agnostic to safety. We find no evidence for a distinct safety subspace. Updates that influence safety, positively or negatively, do not share dominant directions. Any shared structure reflects general learning capacity rather than safety-specific behavior. As such, geometric separation of alignment remains elusive, and linear subspace methods cannot cleanly isolate safety in parameter space.
6 Do Safety Subspaces Exist In Representation Space?
Figure 5: Average Mode Subspace Overlap (MSO) across layers in the 65–90% depth range for pairwise comparisons of activations from Useful (U) and multiple Harmful (H1, H2) prompt sets.
So far, our analysis has focused on the weight space, probing whether certain update directions correspond to safety-related behavior. Across experiments, we found no evidence for distinct subspaces encoding safety at the parameter level. However, safety may instead manifest through how inputs interact with the model, specifically, through the regions of representation space they activate. This motivates a final question: do safety-relevant inputs elicit distinct activation patterns, even if their corresponding weight updates overlap? While weight updates may distribute energy broadly, inputs could selectively activate specific directions. This perspective also offers a possible explanation for earlier results: even low-energy projections onto alignment directions produced strong behavioral effects, likely because inputs activated those directions disproportionately.
Experimental Setup.
We compare internal activations induced by different prompt categories. Specifically, we pass useful (benign) prompts from the MATH dataset [17] and harmful prompts from BeaverTails (test set) and ToxiGen [14] through three models: the aligned model, the useful fine-tuned model, and the contaminated fine-tuned model. For each prompt, we record the hidden state of the last generated token at each transformer layer $\ell\in{0,\dots,L}$. At each layer, these hidden states are stacked into activation matrices of shape $\mathbb{R}^{n\times d}$, where $d$ is the model’s hidden size and $n$ is the number of prompts (5000 for each dataset). We compute MSO (see Section 2) between activation matrices corresponding to the prompts from different datasets, sweeping over energy thresholds $\eta$. Lower values of $\eta$ capture high-energy activation modes, while higher values approximate full-rank comparisons. We plot MSO curves alongside the random-subspace baseline $\max(k_{\text{Useful}},k_{\text{Harmful}})/d$, and report averages over layers in the 65–90 % depth percentile.
Results: Representation Subspaces Overlap Across Tasks.
Figure 5 reports MSO values across all pairs of prompt categories. Useful and harmful prompts consistently exhibit overlap above the random baseline, indicating activation of shared high-energy subspaces in representation space. Notably, the overlap between the two harmful prompt sets is not consistently higher than their overlap with helpful prompts; in some cases, the useful–harmful overlap is greater than the harmful–harmful one. The degree of overlap also varies across model configurations. Some models show strong alignment even in the top subspaces, while others exhibit more gradual increases in overlap, becoming significant only at higher energy thresholds. This variability suggests that representational similarity is influenced more by model-specific factors than by the safety content of the prompts alone. Results on more models are provided in Figure 6 (Appendix D).
Implications: Shared Subspaces Drive Behavior, Not Safety.
These observations suggest that while all prompt types activate shared subspaces more than expected by chance, there is no evidence of a distinct safety-violating subspace. If such a subspace existed, activations from harmful prompts would consistently exhibit greater mutual overlap than with useful prompts, which is not the case. Instead, the results indicate that prompts with differing safety implications are processed through broadly overlapping representations. This supports our earlier hypothesis: the directions most responsible for driving behavior reflect general-purpose representational subspaces rather than safety-specific ones. These directions are activated across tasks and prompt types, implying that LLMs do not internally separate “safe” and “unsafe” activation modes, but instead rely on shared, high-impact subspaces. We find no evidence of a distinct safety subspace in representation space. Useful and harmful prompts show substantial overlap, even across prompt sets with very different behavioral consequences. Combined with our findings in weight space, these results suggest that both aligned and harmful behaviors emerge from shared representational mechanisms rather than separable subspaces.
7 Conclusion
This work set out to investigate how safety alignment is encoded in LLMs, and whether it can be isolated through geometric structure in weight or activation space. Motivated by the challenge of preserving alignment under continued fine-tuning, particularly in adversarial or contaminated settings, we conducted a systematic study across four experiments and five open-source LLMs, examining both parameter updates and internal representations. Our findings challenge the common assumption that alignment corresponds to safety-specific subspaces. Subspaces with high behavioral impact are not unique to alignment; they enhance both utility and harmfulness, and their removal degrades both. This indicates that these directions reflect general-purpose learning rather than safety alone. Moreover, harmful and helpful prompts activate overlapping regions of representation space, offering no evidence for distinct “safety activation” geometry. Together, these results suggest that safety alignment is not cleanly separable in geometric terms. While this complicates subspace-based defenses, it also highlights the potential of high-impact directions, if appropriately constrained, for guiding both safe fine-tuning and activation-level control. More broadly, our work calls for rethinking geometric assumptions in interpretability and alignment, and for developing methods that engage with the entangled nature of learned representations.
8 Limitations
Our study focuses on linear subspaces, providing a principled first step toward understanding the geometric structure of safety alignment. While we do not explore non-linear representations, our framework could be extended in future work to capture richer geometric phenomena. Due to computational constraints, we conduct experiments on small and mid-sized LLMs with publicly available base and instruction-tuned variants. These models offer a controlled and interpretable setting, though generalization to larger architectures and alternative alignment methods remains an important direction for future investigation.
Acknowledgements and Disclosure of Funding
This work received support from Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) and was partially funded by the ADIA Lab Fellowship.
Appendix A Related Work
Safety Alignment and Task-Specific Fine-Tuning in LLMs.
Large Language Models (LLMs) do not inherently follow instructions and often exhibit socially undesirable behaviors. To address this, various post-training methods, instruction-tuning and reinforcement learning from human feedback, are applied to align base LLMs with human values and improve their instruction-following capabilities [37, 46, 55, 44]. However, studies have shown that fine-tuning these aligned models on harmful data can undo this alignment, restoring their original, socially unacceptable behaviors [59]. This unalignment phenomenon has been demonstrated in both open-source models [60, 29] and proprietary models [42, 4, 64] via publicly available fine-tuning APIs, thereby exposing a new attack surface [23, 9, 27]. Moreover, even fine-tuning on benign downstream tasks can degrade alignment [16, 15].
Defense Methods.
To safeguard aligned LLMs against unalignment during fine-tuning, defenses have been proposed at three stages of the pipeline: the alignment stage, the fine-tuning stage, and the post-processing stage. The effectiveness of these defense methods is evaluated using downstream model utility and harmfulness [20].
Alignment Stage Defenses.
Alignment stage defenses update the initial instruction-tuning process to ensure that downstream fine-tuning cannot easily overwrite the model’s safety behavior. One approach augments the alignment loss, making harmful representations harder to recover during fine-tuning updates [45]. Another line of work relies on safety-oriented data curation to preserve alignment under downstream fine-tuning[33]. Adversarial and meta-learning techniques have also been combined to develop tamper-resistant methods that prevent harmfulness while maintaining task performance [48]. A separate strategy introduces a regularization term to the alignment loss, which has been shown to preserve safety after fine-tuning [22]. Perturbing safety-critical layers during instruction-tuning has also been shown to protect alignment [32]. Additional work traces unalignment to excessive dependence on maximum-likelihood training, motivating an integrity preserving variant of this method [6]. A study on “shallow alignment” also shows that instruction-tuning influences only the first few output tokens, whereas deeper alignment improves robustness [41].
Fine-Tuning Stage Defenses.
Fine-tuning stage defenses modify the fine-tuning process to ensure that the model’s alignment is preserved after update. One class of defenses focuses on data curation, augmenting the fine-tuning dataset to maintain alignment after update [5, 13]. Another approach uses safety examples prefixed with a secret prompt, which act as backdoor triggers to reactivate safe behavior after fine-tuning [52]. A data ranking based strategy has also been proposed, where low-quality data is down-ranked and high-quality data is up-ranked to better preserve safety [47]. It has also been shown that prompt templates play an important role; removing the safety prompt during fine-tuning and reintroducing it at inference time can maintain alignment [35].
Optimization based defenses are another type of fine-tuning stage defenses. One line of work splits fine-tuning into an alignment phase and a utility phase, safeguarding both safety and task performance [21]. Another approach combines safety and helpfulness objectives into a single loss [65].
Parameter level methods can also be used to preserve safety. One strategy identifies safety neurons and updates only those parameters [66]. Another approach involves localizing safety layers and freezing their gradients, which has been shown to prevent unalignment [31]. Another line of work explores constraining parameter changes to directions orthogonal to existing safety features, showing that this method preserves alignment [30]. It has also been shown that harmful data can be filtered by matching fine-tuning embeddings against the top-k singular vectors of an activation matrix generated using a harmful dataset [7].
Post-Processing Stage Defenses.
Post-processing stage defenses adjust the fine-tuned model to restore alignment and preserve usefulness. One approach adds a safety vector, defined as the difference between aligned and unaligned weights, to the fine-tuned parameters to regain safe behavior [3]. Another line of work projects the fine-tuning update onto the alignment vector when their similarity drops below a threshold, or selectively merges layers from the fine-tuned and aligned models under the same criterion to achieve a similar effect [18, 10]. A third strategy removes parameters identified as harmful after fine-tuning to restore alignment [19]. It has been shown that safety directions in attention-head activations can also be located and used for targeted intervention [67] to realign the fine-tuned model. Another method detects update parameters whose signs contradict the original alignment and removes them [57]. Additional work restores safety-critical neurons [61], fuses aligned and fine-tuned models [62], or adds an optimized post-hoc perturbation to recover alignment [53].
Safety Mechanisms in Fine-Tuned and Aligned LLMs.
Recent studies have examined how LLMs express safety over neurons, layers, and activations. One study finds that safety related information is language agnostic, identifies parameters whose modification affects alignment, and shows that freezing these parameters during fine-tuning does not ensure safety [40]. Another line of work locates sparse regions in parameter space whose removal weakens alignment, and likewise observes that freezing these regions alone is insufficient to maintain model alignment [54]. A separate analysis maps a safety basin in weight space, noting that random perturbations inside the basin leave safety intact, whereas fine-tuning moves weights outside it [39]. Finally, work on the activation residual stream isolates a refusal direction, removing this direction prevents refusal to harmful prompts, while adding it triggers refusal to benign ones [2].
Appendix B Do Alignment Subspaces Encode Safety?
We provide additional results in Table 3 to support the analysis presented in Section 3.
Table 3: Parallel projection-based update schemes across varying SVD fractions. We report the utility for models fine-tuned on Full Useful data, and harmfulness for models fine-tuned on Full Harmful.
Model Method Utility ($\uparrow$) SVD Fractions Harmful Score ($\downarrow$) SVD Fractions
0.01 0.25 0.50 0.75 0.99 0.01 0.25 0.50 0.75 0.99
Qwen-2.5 1.5B Base $0.21$ $3.27$
Aligned $0.47$ $1.55$
Fine-Tuned $0.61$ $2.09$
Top-K $0.50$ $0.53$ $0.55$ $0.57$ $0.58$ $1.62$ $1.80$ $1.92$ $1.90$ $1.97$
Random-K $0.49$ $0.50$ $0.53$ $0.56$ $0.58$ $1.55$ $1.66$ $1.78$ $1.92$ $2.00$
Random $0.49$ $0.50$ $0.53$ $0.53$ $0.56$ $1.56$ $1.65$ $1.74$ $1.83$ $1.95$
Llama-3.2 1B Base $0.03$ $4.13$
Aligned $0.13$ $2.80$
Fine-Tuned $0.36$ $4.07$
Top-K $0.14$ $0.21$ $0.25$ $0.30$ $0.34$ $2.89$ $3.29$ $3.51$ $3.66$ $3.84$
Random-K $0.13$ $0.16$ $0.23$ $0.29$ $0.34$ $2.83$ $3.11$ $3.37$ $3.55$ $3.84$
Random $0.13$ $0.17$ $0.22$ $0.29$ $0.34$ $2.81$ $3.05$ $3.34$ $3.56$ $3.83$
Qwen-2.5 3B Base $0.44$ $2.53$
Aligned $0.61$ $1.47$
Fine-Tuned $0.72$ $2.16$
Top-K $0.63$ $0.64$ $0.65$ $0.68$ $0.69$ $1.48$ $1.71$ $1.81$ $1.91$ $1.92$
Random-K $0.62$ $0.63$ $0.64$ $0.65$ $0.69$ $1.44$ $1.55$ $1.62$ $1.74$ $1.91$
Random $0.62$ $0.63$ $0.64$ $0.65$ $0.68$ $1.44$ $1.50$ $1.66$ $1.75$ $1.83$
Qwen-2.5 7B Base $0.69$ $1.90$
Aligned $0.74$ $1.30$
Fine-Tuned $0.81$ $2.12$
Top-K $0.72$ $0.74$ $0.76$ $0.77$ $0.77$ $1.34$ $1.56$ $1.66$ $1.76$ $1.84$
Random-K $0.73$ $0.75$ $0.74$ $0.75$ $0.77$ $1.34$ $1.44$ $1.53$ $1.64$ $1.84$
Random $0.74$ $0.75$ $0.75$ $0.76$ $0.76$ $1.33$ $1.40$ $1.48$ $1.56$ $1.75$
Llama-2 7B Base $0.05$ $4.27$
Aligned $0.20$ $1.74$
Fine-Tuned $0.30$ $3.41$
Top-K $0.21$ $0.24$ $0.26$ $0.28$ $0.29$ $1.81$ $2.34$ $2.61$ $2.90$ $3.15$
Random-K $0.20$ $0.23$ $0.25$ $0.28$ $0.29$ $1.74$ $1.91$ $2.09$ $2.63$ $3.13$
Random $0.20$ $0.23$ $0.25$ $0.28$ $0.28$ $1.77$ $1.91$ $2.15$ $2.57$ $3.03$
Appendix C Can Harmful Subspaces Be Removed?
Table 4 presents supplementary results that further substantiate the findings discussed in Section 4.
Table 4: Parallel projection-based update schemes across varying SVD fractions. We report the utility and harmfulness for models fine-tuned on Contaminated data.
Model Method Utility ($\uparrow$) emphSVD Fractions Harmful Score ($\downarrow$) emphSVD Fractions
0.01 0.25 0.50 0.75 0.99 0.01 0.25 0.50 0.75 0.99
Qwen-2.5 1.5B Base $0.21$ $3.27$
Aligned $0.47$ $1.55$
FT $0.60$ $2.16$
Top-K $0.50$ $0.53$ $0.52$ $0.55$ $0.56$ $1.59$ $1.65$ $1.79$ $1.91$ $1.92$
Random-K $0.49$ $0.52$ $0.53$ $0.55$ $0.55$ $1.56$ $1.62$ $1.63$ $1.87$ $1.92$
Random $0.49$ $0.50$ $0.52$ $0.52$ $0.54$ $1.58$ $1.64$ $1.68$ $1.74$ $1.92$
Llama-3.2 1B Base $0.026$ $4.13$
Aligned $0.13$ $2.80$
FT $0.37$ $3.60$
Top-K $0.14$ $0.20$ $0.25$ $0.29$ $0.33$ $2.84$ $2.90$ $3.05$ $3.36$ $3.45$
Random-K $0.13$ $0.16$ $0.22$ $0.29$ $0.33$ $2.81$ $2.90$ $3.03$ $3.19$ $3.45$
Random $0.13$ $0.16$ $0.22$ $0.28$ $0.33$ $2.84$ $2.90$ $3.19$ $3.19$ $3.45$
Qwen-2.5 3B Base $0.44$ $2.53$
Aligned $0.61$ $1.47$
FT $0.73$ $1.99$
Top-K $0.62$ $0.63$ $0.65$ $0.68$ $0.69$ $1.49$ $1.58$ $1.69$ $1.76$ $1.83$
Random-K $0.62$ $0.64$ $0.64$ $0.66$ $0.69$ $1.45$ $1.55$ $1.62$ $1.65$ $1.91$
Random $0.62$ $0.63$ $0.64$ $0.65$ $0.68$ $1.45$ $1.50$ $1.57$ $1.75$ $1.83$
Qwen-2.5 7B Base $0.69$ $1.90$
Aligned $0.74$ $1.30$
FT $0.81$ $1.96$
Top-K $0.74$ $0.75$ $0.75$ $0.75$ $0.78$ $1.30$ $1.55$ $1.60$ $1.68$ $1.67$
Random-K $0.74$ $0.75$ $0.76$ $0.75$ $0.78$ $1.35$ $1.41$ $1.46$ $1.59$ $1.67$
Random $0.74$ $0.75$ $0.75$ $0.75$ $0.78$ $1.34$ $1.40$ $1.48$ $1.56$ $1.63$
Llama-2 7B Base $0.053$ $4.27$
Aligned $0.20$ $1.74$
FT $0.30$ $3.08$
Top-K $0.21$ $0.23$ $0.25$ $0.27$ $0.28$ $1.77$ $1.91$ $2.15$ $2.38$ $2.74$
Random-K $0.20$ $0.23$ $0.26$ $0.28$ $0.28$ $1.74$ $1.91$ $2.09$ $2.38$ $2.79$
Random $0.20$ $0.23$ $0.25$ $0.27$ $0.28$ $1.77$ $1.91$ $2.15$ $2.38$ $2.74$
Appendix D Do Safety Subspaces Exist in Representation Space?
To complement the discussion in Section 5, we include extended results in Figure 6.
Figure 6: Average Mode Subspace Overlap (MSO) across layers in the 65–90% depth range for pairwise comparisons of activations from Useful (U) and multiple Harmful (H1, H2) prompt sets.
Appendix E Experimental Details
We implemented all experiments using PyTorch [38] and the HuggingFace Transformers library [56]. We ran all experiments on a single NVIDIA A6000 GPU (48 GB). To save memory, all base models are initalized in torch.bfloat16 precision. All models are trained using the AdamW optimizer [34]. Detailed hyperparameter configurations for full fine-tuning of each model are presented in Table 5.
Table 5: Hyperparameter settings for fine-tuning the various models.
Optimizer AdamW
Batch size 1
Max. Seq. Len 512
Grad Acc. Steps 32
Epochs 1
Learning Rate $1\times 10^{-5}$
LR Scheduler Cosine
Warmup Ratio 0.02
Appendix F Dataset Details
We use the MetaMathQA dataset [63] for fine-tuning, which reformulates existing math problems from alternative perspectives without introducing new content. To evaluate performance, we rely on the GSM8K benchmark [8], a dataset of elementary-level math questions that require multi-step reasoning. Models are assessed based solely on the correctness of the final numerical answer. For our activation-based analysis, we sample prompts from the MATH dataset [17], which contains challenging, competition-style arithmetic problems.
BeaverTails [26] is a valuable dataset for studying safety by independently annotating question–answer pairs for both helpfulness and harmlessness. We use the training set to fine-tune models in both harmful and contaminated settings, and draw prompts from the test split for our activation-based experiments.
AdvBench [68] consists of 500 prompts designed to elicit a wide range of harmful behaviors, including profanity, threats, misinformation, discrimination, cybercrime, and other forms of dangerous or illegal content framed as instructions. We use this benchmark to quantify model harmfulness: higher success in responding to these prompts indicates greater unsafe behavior.
ToxiGen [14] is a large-scale dataset composed of both toxic and non-toxic statements. We use a subset of its prompts to analyze model activations in response to harmful content.